# Intrinsic nonlinear Hall effect beyond Bloch geometry

### v1

@@ Abstract
The theory of the intrinsic Hall effect, both linear and nonlinear, is rooted in a geometry which is defined in the Bloch-vector parameter space; the formal expressions are mostly derived from semiclassical concepts. When disorder and interaction are considered there is no Bloch vector to speak of; one needs a more general quantum geometry, defined in a different parameter space. Such higher-level geometrical formulation of the intrinsic Hall effect provides very compact expressions, which have the additional virtue—in the Bloch special case—of yielding the known results in a straightforward way: the logic is not concealed by the algebra.
Introduction.–The second-order dc conductivity tensor is comprised of three terms which—in a semiclassical treatment—scale as the zeroth, first, and second power of the transport lifetime $\tau$. The first two terms are of the Hall kind, i.e. the induced current is normal to the electric field, and their main entries are quantum-geometrical entities: Berry connections and curvatures in Bloch space [1, 2, 3, 4, 5]. The main focus of the present work is on the $\tau^{0}$ term, which—to the best of the author’s knowledge—was first discovered in Ref. [1], where it was attributed to the field-induced “positional shift” of Bloch electrons; it will be indicated as $\sigma^{(\rm ps)}_{\alpha\beta\gamma}$ in the following. This term has received much attention in recent times [6, 7, 8, 9, 10, 11].
I adopt here an exact quantum-mechanical framework, where by “exact” I mean that I address a system of interacting electrons and static classical nuclei, possibly in a disordered configuration. No relaxation time $\tau$ can enter the theory: in fact the concept itself does not make sense beyond the semiclassical theory. In the present framework the response functions are causal but nondissipative: therefore the induced current grows with some powers of time. At first order the intrinsic Hall current is $t^{0}$, and the Drude current is $t^{1}$ (the electrons undergo free acceleration). At second order one expects three terms, namely $t^{0}$, $t^{1}$, and $t^{2}$, in one-to-one correspondence to the three semiclassical terms; while for two of them a many-body formulation exists [12, 13], the $t^{0}$ term remained elusive so far. Its expression is provided here; for the sake of completeness the other terms are also addressed.
@@
Exact theory.–The flux-dependent many body Hamiltonian has been introduced in 1964 by Kohn [21]; when modified in order to account for the absence of time-reversal (T) symmetry it reads
@@
The linear conductivity obtains by setting ${\mbox{$\kappa$}}=0$; switching from the $t$-domain to the $\omega$-domain and enforcing causality the second term in Eq. (4) yields the familiar linear Drude conductivity
$\sigma^{(\rm Drude)}_{\alpha\beta}(\omega)=\frac{D_{\alpha\beta}}{\pi}\frac{i}{\omega+i\eta}=D_{\alpha\beta}\left[\delta(\omega)+\frac{i}{\pi\omega}\right],$ (5)
where $\eta=0^{+}$ and
$D_{\alpha\beta}=\frac{\pi e^{2}}{\hbar^{2}L^{d}}\frac{\partial^{2}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}}$ (6)
evaluated at ${\mbox{$\kappa$}}=0$ is known as the Drude weight [21, 22, 23].
@@
$\partial_{t}D_{\alpha\beta}$ $=$ $\frac{\pi e^{2}}{\hbar^{2}L^{d}}\frac{\partial^{3}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}\partial\kappa_{\gamma}}\dot{\kappa}_{\gamma}.$ (7)
@@
- $\frac{\pi e^{3}t^{2}}{\hbar^{3}L^{d}}\frac{\partial^{3}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}\partial\kappa_{\gamma}},$
@@
$\lim_{\omega\rightarrow 0}\frac{\mbox{Im }\sigma^{(\rm regular)}_{\alpha\gamma}(\omega)}{\omega}=\frac{e}{L^{d}}\Omega(\kappa_{\alpha},{\cal E}_{\gamma}).$ (15)

### v2

@@ Abstract
The theory of the intrinsic Hall effect, both linear and nonlinear, is rooted in a geometry which is defined in the Bloch-vector parameter space; the formal expressions are mostly derived from semiclassical concepts. When disorder and interaction are considered there is no Bloch vector to speak of; one needs a more general quantum geometry, defined in a different parameter space. The nonlinear Hall effect is a fundamental geometric response of the many-body ground state, not a band-structure peculiarity. The higher-level geometrical formulation of the intrinsic Hall effect provides very compact expressions, which have the additional virtue—in the Bloch special case—of yielding the known results in a straightforward way: the logic is not concealed by the algebra.
Introduction.–The second-order dc conductivity tensor is comprised of three terms which—in a semiclassical treatment—scale as the zeroth, first, and second power of the transport lifetime $\tau$. The first two terms are of the Hall kind, i.e. the induced current is normal to the electric field, and their main entries are quantum-geometrical entities: Berry connections and curvatures in Bloch space [1, 2, 3, 4, 5]. The main focus of the present work is on the $\tau^{0}$ term, which—to the best of the author’s knowledge—was first discovered in Ref. [1], where it was attributed to the field-induced “positional shift” of Bloch electrons; it will be indicated as $\sigma^{(\rm ps)}_{\alpha\beta\gamma}$ in the following. This term has received much attention in recent times [6, 7, 8, 9, 10, 11]; it is shown here that it is not a peculiarity of Bloch electrons: it is a fundamental geometric response of the many-body ground state.
I adopt here an exact quantum-mechanical framework, where by “exact” I mean that I address a nonrelativistic system of interacting electrons and static classical nuclei, possibly in a disordered configuration. No relaxation time $\tau$ can enter the theory: in fact the concept itself does not make sense beyond the semiclassical theory. In the present framework the response functions are causal but nondissipative: therefore the induced current grows with some powers of time. At first order the intrinsic Hall current is $t^{0}$, and the Drude current is $t^{1}$ (the electrons undergo free acceleration). At second order one expects three terms, namely $t^{0}$, $t^{1}$, and $t^{2}$, in one-to-one correspondence to the three semiclassical terms, having the same symmetry selection rules: the $t^{0}$ and $t^{2}$ terms require breaking of time-reversal (T) symmetry, besides breaking of inversion symmetry [5]. While for the $t^{1}$ and $t^{2}$ terms a many-body formulation exists [12, 13], the $t^{0}$ term remained elusive so far. Its expression is provided here; for the sake of completeness the other terms are also addressed.
@@
Exact theory.–The flux-dependent many body Hamiltonian has been introduced in 1964 by Kohn [21]; the most general nonrelativistic Hamiltonian, also accounting for the absence of T-symmetry, reads
@@
The linear conductivity obtains by setting ${\mbox{$\kappa$}}=0$: the two terms in Eq. (4) are $t^{0}$ (transverse) and $t^{1}$ (longitudinal), respectively. The longitudinal term in Eq. (4) is rewritten as
$\partial_{t}j(t)=\frac{D_{\alpha\beta}}{\pi}{\cal E}_{\beta},\quad D_{\alpha\beta}=\frac{\pi e^{2}}{\hbar^{2}L^{d}}\frac{\partial^{2}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}},$ (5)
where $D_{\alpha\beta}$ at ${\mbox{$\kappa$}}=0$ is known as the Drude weight [21, 22, 23]; the Fourier transform of Eq. (5) is
$i\omega j_{\alpha}(\omega)=\frac{D_{\alpha\beta}}{\pi}{\cal E}_{\beta}(\omega),$ (6)
whose causal inversion yields
$j_{\alpha}(\omega)=\frac{D_{\alpha\beta}}{\pi}\frac{i}{\omega+i\eta}{\cal E}_{\beta}(\omega),\quad\eta=0^{+}.$ (7)
The famous Kohn’s expression [21] for the Drude term in longitudinal conductivity is
$\sigma^{(\rm Drude)}_{\alpha\beta}(\omega)=\frac{D_{\alpha\beta}}{\pi}\frac{i}{\omega+i\eta}=D_{\alpha\beta}\left[\delta(\omega)+\frac{i}{\pi\omega}\right].$ (8)
@@
$\partial_{t}D_{\alpha\beta}$ $=$ $\frac{e^{2}}{\hbar^{2}L^{d}}\frac{\partial^{3}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}\partial\kappa_{\gamma}}\dot{\kappa}_{\gamma}.$ (9)
@@
- $\frac{e^{3}t^{2}}{\hbar^{3}L^{d}}\frac{\partial^{3}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}\partial\kappa_{\gamma}},$
@@
$\lim_{\omega\rightarrow 0}\frac{\mbox{Im }\sigma^{(\rm regular)}_{\alpha\gamma}(\omega)}{\omega}=\frac{e}{L^{d}}\Omega({\cal E}_{\gamma},\kappa_{\alpha}).$ (17)